# Generative agent-based modeling with actions grounded in physical, social, or digital space using Concordia

### 06 Dec 2023 v1

@@ 2.1 Generative agents
A Concordia agent has both a long-term memory and a working memory. Let the long-term memory be a set of strings $\mathbf{m}$ that records everything remembered or currently experienced by the agent. The working memory is $\mathbf{z}=\{z^{i}\}_{i}$ is composed of the states of individual components (Figure 2). A component $i$ has a state $z^{i}$, which is statement in natural language—e.g. “Alice is at work”. The components update their states by querying the memory (which contains the incoming observations) and using LLM for summarising and reasoning. Components can also condition their update on the current state of other components. For example, the planing component can update its state if an incoming observation invalidates the current plan, conditioned on the state of the ‘goal’ component. Components can as well have internal logic that is programmed in using classic programming, for example a hunger component can check how many calories and how long ago did the agent consume and update its state based on that.
We use the same associative memory architecture as in Park et al. (2023) [^3] The idea of simulating a group of generative agents has been explored in a variety of ways in recent work. Our work is focused on on agent-based modeling for science and for evaluation of digital technologies. Another recent line of work has focused instead on the idea of using groups of generative agents to simulate organizations that solve problems like software companies and to thereby try to build a general-purpose problem solving system (Hong et al., 2023; Li et al., 2023b).. We feed the incoming observations immediately into the agents memory, so they are available during when components update [^4] For convenience, we also allow the components to subscribe to the observation stream explicitly..
@@
Formally, the agent is defined as a two step sampling process, using a LLM $p$ (see Figure 3 for illustration). In the action step, the agent samples it’s activity $a_{t}$, given the state of components $\mathbf{z}_{t}=\{z_{t}^{i}\}_{i}$:
@@
In the second step the agent samples it’s state $\mathbf{z}$, given the agents memory $\mathbf{m}_{t}$ up to the present time:
@@
Notice how eq.1 and eq.2 are not fundamentally different. What makes a difference between an agent output and a component is that the output of the former is interpreted as an action in the environment. In eq.1 we also don’t explicitly condition on the memory to point out the architectural decision, where components mediate between a long-term memory and the agents working memory. Otherwise, we can think of agent as a special component and of components as sub-agents.
@@ 2.2 Generative environments
Here we explicitly condition on the action attempted by the agent, although it could be subsumed into the components (like observation in eq.1). This is to highlight that the GM generates an event statement $e_{t}$ in response to every action of any agent, while the agent might take in several observations before it acts (or none at all). After adding the event statement $e_{t}$ to its memory the GM can update it’s components using the same eq. 2 as the agent. It can then emit observations $\mathbf{o}^{i}_{t}$ for player $i$ using the following equation:
@@ 3.1 Neuroscience interpretation of the generative agent architecture
To implement routine behavior, an agent could continually rehearse its routine in working memory, but that would impair its ability to use working memory for other purposes on other tasks since its working memory is limited in capacity. So instead of continually rehearsing routines in working memory, we may instead assume that they are often stored elsewhere and then retrieved when needed (i.e. from long-term memory).
@@ 3.3.1 Concordia agents are not reinforcement learners
Generative view of agency presented in this paper contrasts with the classic Reinforcement Learning (RL) view as summarized in the “Reward is enough” thesis of Silver et al. (2021). The orthodox RL view of behaviour is that it is constructed from individual experience and driven by a quantifiable (and externally supplied) reward function reflecting the achievement of goals. To communicate what behaviour is desired of the agent, one has to annotate the agents’ activity with a reward signal, which signals goal achievement. Here we instead follow the social constructionist view of agency expressed in (March and Olsen, 2011), where behaviour is an expression of the agents position in the social context, and what policy the social norms prescribe for the agent in such a position. Answering “what does a person such as I do in a situation such as this?” might require positing a practical goal and achieving it (“make money”, “get famous”), but goals are qualitative, dynamic and context dependent. To specify the behavior you want an agent to produce you need to communicate its social context and the agents position within it.
@@
Notice, that this doesn’t mean that Concordia agents couldn’t, in principle, perform reward maximisation and policy iteration. Brooks et al. (2023) have shown that the ability of LLMs to learn in-context Brown et al. (2020) can be used to perform policy iteration in classic RL environments, as long as they can be represented as text. One could also implement a specialised component that runs a classic RL algorithm for a specific domain or tool use case. The agent could provide supervision to its RL based components via heirarcical RL techniques like feudal RL Dayan and Hinton (1992); Vezhnevets et al. (2017).
@@ 4.3 Sequential social dilemmas experiments in silico
Concordia adds to the toolbox for studying multi-agent problems such as resource management, social dilemmas, commons problems, cooperation, equilibrium selection, and coordination (Leibo et al., 2017, 2021). Previously these problems have either been cast as matrix games or as multi-agent RL (MARL) (Hertz et al., 2023). We see generative agents as the next step in the evolutionary line of ‘homo-economicus’ and ‘homo-RLicus’.
GABM makes it possible to investigate how rules, laws and social norms formulated in language influence, for example, the management of shared resources. Moreover, we can investigate whether the demands of sharing a resource give rise to rules, laws and norms capable of governing that resource (and under what circumstances this works or does not)—i.e. whether rules are emergent, and what the conditions are for their emergence. For example, Hadfield and Weingast (2013) proposed that legal order can emerge without the centralised enforcement in certain demographic and economical circumstances. The demonstrate it using historical examples from gold-rush in California and medieval Iceland. Concordia could be used to simulate those examples and enable further insights into the nature of legal order. For example, we could check whether certain demographic assumptions are necessary by varying the number of agents.
@@ 4.4 Concordia can implement classic and contemporary psychological models
Many influential psychological models have distinguished between more associative and more deliberative processes for decision-making (e.g. Schneider and Shiffrin (1977); Kahneman et al. (2002); Dayan (2009)). Whereas implicit-associative processes learn the regularity of the world slowly for intuitive judgment, the explicit-deliberative processes are thought to be more linguistically mediated and allow for symbolic inference and faster learning in novel situations (Greenwald and Banaji (1995); Wilson et al. (2000)). Because the implicit-associative models are conceptually easy to model within connectionist or neural network frameworks (Smith (2009)), ABMs have been more closely aligned with models that focus on the associative decision making processes, or the parts of the model that are associative and neglecting their more symbolic and deliberative aspects. Many of these more symbolic psychological models take an “arrow and box” approach to theorizing which describe high level processes and transformations of information, and often posit sequential steps of information flow. Now using generative agents like Concordia such symbolic and deliberative aspects of cognition are also easy to capture in computational models.
@@ 4.6 Emergence and multi-scale modeling with Concordia
Demonstrating the emergence of a particular social phenomena from the behaviour of individual agents, which are not explicitly instructed to produce it, is important an important topic in multi-agent research (Walker and Wooldridge, 1995; Leibo et al., 2019, 2021; Axtell et al., 2001). Indeed, the phenomena of intelligence is also hypothesised to be an emergent social phenomena (Duéñez-Guzmán et al., 2023). DeLanda (2011), for example, explores the topic of emergence and simulation across various fields. So far many multi-agent simulation approaches have been limited by being able to only deal with one, fixed scale of the simulation—individual agents. Scaling further is computationally challenging.
Concordia allows modelling systems across multiple scales, where phenomena at each scale constitute a substrate for the emergence of the phenomena on the next scale (Koestler, 1967; DeLanda, 2011). For example, individual agents form a substrate from which social institutions and organisations can arise. Through engaging in exchange of goods and services, the agents can create an economy and, for example, start a bank. Modelling a banking system this way would be, most likely, computationally prohibitive. Since in Concordia the agents (or GM) need not represent individuals, but could be organisations, institutions or even nation states, we could enrich the simulation by adding generative agent versions of other banks and businesses. They will be modelled in courser resolution (not as emergent from individual activities), but could be made accurate. Such simulation would allow to model how certain interventions (e.g. central bank interest rate) would propagate across the macro and micro scales of economic activity.
@@ 5 Future work
Since there is no consensus at present concerning how to interpret results of LLM-based simulations of human populations, the future work will address the critical epistemic question: “by what standard should we judge whether (and in what ways, and under which conditions) the results of in silico experiments are likely to generalize to the real world?”. These are not questions any one group of researchers can answer by themselves; rather these issues must be negotiated by the community as a whole. This is is why we release Concordia early and with only few examples. It is an invitation to the researchers from various fields that are interested in GABM to come onboard and participate in the creation of validating procedures are epistemic norms.
@@
    Adding new environments, including integration with different LLMs to see which are more suitable for constructing GABM (e.g., act “reasonably”, internally consistent, applies common sens)
@@
  5. 5.
    Keyframes—conditioning the agent actions to be consistent with future key action or of narrative. This allow steering the simulation more granularly and addresses an inherent issue that is caused by the fact that there is no guarantee that due to the random nature of GABMs, simulation might diverge completely from the scenario.
@@ A.8 Examples
    Calendar: An illustrative social simulation with 2 players which simulates phone interactions. The two players, Alice and Bob, have a smartphone with a Calendar app. Alice’s goal is to setup a meeting with Bob using the Calendar app on her phone, taking Bob’s schedulde into account when selecting the date/time.
@@
    Riverbend elections: An illustrative social simulation with 5 players which simulates the day of mayoral elections in an imaginary town caller Riverbend. First two players, Alice and Bob, are running for the mayor. The third player, Charlie, is trying to ruin Alices’ reputation with disinformation. The last two players have no specific agenda, apart from voting in the election.
@@
    Magic Beans for sale: An example illustrating how to use the inventory component. Agents can buy and trade magic beans.
@@
    Cyberball: An example which simulates social exclusion experiment Williams et al. (2000) and shows how to use a standard psychology questionnaire.

### 13 Dec 2023 v2

@@ 2.1 Generative agents
A Concordia agent has both a long-term memory and a working memory. Let the long-term memory be a set of strings $\mathbf{m}$ that records everything remembered or currently experienced by the agent. The working memory is $\mathbf{z}=\{z^{i}\}_{i}$ is composed of the states of individual components (Figure 2). A component $i$ has a state $z^{i}$, which is statement in natural language—e.g. “Alice is at work”. The components update their states by querying the memory (which contains the incoming observations) and using LLM for summarising and reasoning. Components can also condition their update on the current state of other components. For example, the planning component can update its state if an incoming observation invalidates the current plan, conditioned on the state of the ‘goal’ component. Components can also have internal logic programmed using classic programming, for example a hunger component can check how many calories an agent consumed and how recently it consumed them, and update its state based on the result.
We use the same associative memory architecture as in Park et al. (2023) [^3] The idea of simulating a group of generative agents has been explored in a variety of ways in recent work. Our work is focused on on agent-based modeling for science and for evaluation of digital technologies. Another recent line of work has focused instead on the idea of using groups of generative agents to simulate organizations that solve problems like software companies and to thereby try to build a general-purpose problem solving system (Hong et al., 2023; Li et al., 2023b).. We feed the incoming observations immediately into the agents memory, to make them available when components update [^4] For convenience, we also allow the components to subscribe to the observation stream explicitly..
@@
Formally, the agent is defined as a two step sampling process, using a LLM $p$ (see Figure 3 for illustration). In the action step, the agent samples its activity $a_{t}$, given the state of components $\mathbf{z}_{t}=\{z_{t}^{i}\}_{i}$:
@@
In the second step the agent samples its state $\mathbf{z}$, given the agents memory $\mathbf{m}_{t}$ up to the present time:
@@
Notice how eq.1 and eq.2 are not fundamentally different. What makes the difference between an agent output and a component is that the output of the former is interpreted by the GM as an action in the environment. In eq.1 we also don’t explicitly condition on the memory to point out the architectural decision, where components mediate between a long-term memory and the agents working memory. Otherwise, we can think of an agent as a special kind of component and of components as sub-agents.
@@ 2.2 Generative environments
Here we explicitly condition on the action attempted by the agent, although it could be subsumed into the components (like observation in eq.1). This is to highlight that the GM generates an event statement $e_{t}$ in response to every action of any agent, while the agent might take in several observations before it acts (or none at all). After adding the event statement $e_{t}$ to its memory the GM can update its components using the same eq. 2 as the agent. It can then emit observations $\mathbf{o}^{i}_{t}$ for player $i$ using the following equation:
@@ 3.1 Neuroscience interpretation of the generative agent architecture
To implement routine behavior, an agent could continually rehearse its routine in working memory, but that would impair its ability to use working memory for other purposes on other tasks since its working memory is limited in capacity (like in Baddeley (1992)). So instead of continually rehearsing routines in working memory, we may instead assume that they are often stored elsewhere and then retrieved when needed (i.e. from long-term memory).
@@ 3.3.1 Concordia agents are not reinforcement learners
Generative view of agency presented in this paper contrasts with the classic Reinforcement Learning (RL) view as summarized in the “Reward is enough” thesis of Silver et al. (2021). The orthodox RL view of behaviour is that it is constructed from individual experience and driven by a quantifiable (and externally supplied) reward function reflecting the achievement of goals. To communicate what behaviour is desired of the agent, one has to annotate the agents’ activity with a reward signal, which signals goal achievement. Here we instead follow the social constructionist view of agency expressed in March and Olsen (2011), where behavior is an expression of the agent’s position in the social context, and what policy the social norms prescribe for the agent in such a position. Answering “what does a person such as I do in a situation such as this?” might require positing a practical goal and achieving it (“make money”, “get famous”), but goals are qualitative, dynamic and context dependent. To specify the behavior you want an agent to produce you need to communicate its social context and the agents position within it.
@@
Notice, that this doesn’t mean that Concordia agents couldn’t, in principle, perform reward maximisation and policy iteration. Brooks et al. (2023) have shown that the ability of LLMs to learn in-context (Brown et al., 2020) can be used to perform policy iteration in classic RL environments, as long as they can be represented as text. One could also implement a specialised component that runs a classic RL algorithm for a specific domain or tool use case. The agent could provide supervision to its RL based components via hierarchical RL techniques like feudal RL (Dayan and Hinton, 1992; Vezhnevets et al., 2017).
@@ 4.3 Sequential social dilemmas experiments in silico
Concordia adds to the toolbox for studying multi-agent problems such as resource management, social dilemmas, commons problems, cooperation, equilibrium selection, and coordination (Leibo et al., 2017, 2021). Previously these problems have either been cast as matrix games or as multi-agent RL (MARL) (Hertz et al., 2023). Now it is clear that many researchers, including us, see that an LLM-based approach is possible and will have many advantages, as evidenced by the fact that quite a few frameworks for social modeling with LLMs appeared this year (Wu et al., 2023; Zhou et al., 2023; Kaiya et al., 2023). We see generative agents as the next step in the evolutionary line of “model animals” after ‘Homo-economicus’ and ‘Homo-RLicus’.
Generative agent-based modeling makes it possible to investigate how rules, laws and social norms formulated in language influence, for example, the management of shared resources (e.g. Yocum et al. (2023)). With Concordia we will be able to investigate whether the demands of sharing a resource give rise to rules, laws and norms capable of governing that resource (and under what circumstances this works or does not)—i.e. whether rules are emergent, and what the conditions are for their emergence. For example, Hadfield and Weingast (2013) proposed that legal order can emerge without centralised enforcement in certain circumstances. They demonstrate this using historical examples from gold-rush in California and medieval Iceland. Concordia could be used to simulate those examples and enable further insights into the nature of legal order. For example, we could check whether certain demographic assumptions are necessary by varying the number of agents.
@@ 4.4 Concordia can implement classic and contemporary psychological models
Many influential psychological models have distinguished between more associative and more deliberative processes for decision-making (e.g. Schneider and Shiffrin (1977); Kahneman et al. (2002); Dayan (2009)). Whereas implicit-associative processes learn the regularity of the world slowly for intuitive judgment, the explicit-deliberative processes are thought to be more linguistically mediated and allow for symbolic inference and faster learning in novel situations (Greenwald and Banaji (1995); Wilson et al. (2000)). Because the implicit-associative models are conceptually easy to model within connectionist or neural network frameworks (Smith (2009)), many ABMs have been more closely aligned with models of individual decision making that focus on its associative processes or the associative parts of complex models, and have neglected their more symbolic and deliberative aspects. Many of these more symbolic psychological models take an “arrow and box” approach to theorizing which describe high level processes and transformations of information, and often posit sequential steps of information flow. Now using generative agents like Concordia such symbolic and deliberative aspects of cognition are also easy to capture in computational models.
@@ 4.6 Emergence and multi-scale modeling with Concordia
Demonstrating the emergence of a particular social phenomena from the behaviour of individual agents, which are not explicitly instructed to produce it, is important an important topic in multi-agent research (Walker and Wooldridge, 1995; Leibo et al., 2019, 2021; Axtell et al., 2001). Indeed, much of what is distinctive about human intelligence is hypothesised to be an emergent social phenomena involving multi-scale interactions (Henrich, 2016; Wilson et al., 2013). DeLanda (2011), for example, explores the topic of emergence and simulation across various fields. While the wider ABM field has studied multi-scale models (Tesfatsion, 2023), the approaches based on deep reinforcement learning have been limited by being able to only deal with one fixed scale of the simulation: individual agents (e.g. Johanson et al. (2022); Zheng et al. (2022)), and scaling deep RL to large numbers of agents would be computationally difficult.
Concordia allows modeling systems across multiple scales, where phenomena at each scale constitute a substrate for the emergence of the phenomena on the next scale (Koestler, 1967; DeLanda, 2011; Duéñez-Guzmán et al., 2023). For example, individual agents form a substrate from which social institutions and organisations can arise. Through engaging in exchange of goods and services, the agents can create an economy and, for example, start a bank. Modelling a banking system this way would be, most likely, computationally prohibitive. Since in Concordia the agents (or GM) need not represent individuals, but could be organisations, institutions or even nation states, we could enrich simulations by adding generative agent versions of other entities such as banks and businesses. They could be modeled with coarser resolution, not just as emerging from the activities of individual agents, but could be made accurate for instance by incorporating precise models of how they operate. Such simulations could be used to model how interventions (e.g. a central bank interest rate decision) propagate across macro and micro scales of economic activity.
@@ 5 Future work
Since there is no consensus at present concerning how to interpret results of LLM-based simulations of human populations, the future work will address the critical epistemic question: “by what standard should we judge whether (and in what ways, and under which conditions) the results of in silico experiments are likely to generalize to the real world?”. These are not questions any one group of researchers can answer by themselves; rather these issues must be negotiated by the community as a whole. This is is why we release Concordia early and with only few examples. It is an invitation to the researchers from various fields that are interested in GABM to come onboard and participate in the creation of validating procedures, best practices, and epistemic norms.
We plan to add the following over the coming months:
@@
    New example environments
@@
    Integration with different LLMs to see which are more suitable for constructing GABMs (e.g., they act “reasonably”, are internally consistent, apply common sense, etc).
  3. 3.
@@
  6. 6.
    Keyframes—conditioning the agent actions to be consistent with future key action or of narrative. This allow steering the simulation more granularly and addresses an inherent issue that is caused by the fact that there is no guarantee that due to the stochastic nature of GABMs, ongoing simulations might diverge from their intended topic.
@@ A.8 Examples
    Calendar: An illustrative social simulation with 2 players which simulates phone interactions. The two players, Alice and Bob, have a smartphone with a Calendar app. Alice’s goal is to setup a meeting with Bob using the Calendar app on her phone, taking Bob’s schedule into account when selecting the date/time.
@@
    Riverbend elections: An illustrative social simulation with 5 players which simulates the day of mayoral elections in an imaginary town caller Riverbend. First two players, Alice and Bob, are running for the mayor. The third player, Charlie, is trying to ruin Alice’s reputation with disinformation. The last two players have no specific agenda, apart from voting in the election.
@@
    Magic Beans for sale: An example illustrating how to use the inventory component. Agents can buy and trade beans for money.
@@
    Cyberball: An example which simulates social exclusion using a GABM version of a standard social psychology paradigm (Williams et al., 2000) and shows how to use standard psychology questionnaires.